OpenStudy (anonymous):

What value of x makes the equation true? –(9 + (–17)) = –9 + x x =

OpenStudy (anonymous):

Lol one more

OpenStudy (anonymous):

–(9 + (–17)) = –9 + x -9 + 17 = -9 + x; x = -9 + 9 + 17; x = 17

OpenStudy (anonymous):

i hope that helps

OpenStudy (anonymous):

yes thx

OpenStudy (anonymous):

u welcome

OpenStudy (anonymous):

80 percent thx i missed one that i answered nut thx for the help 4/5

OpenStudy (anonymous):

you welcome